1/w+14 + 1/3w-8 = -50/3w^2+34w-112
Answer by stanbon(75887)

1/(w+14) + 1/(3w-8) = -50/(3w^2+34w-112)
---------
Note: 3w^2 + 34w-112 = (w+14)(3w-8)
------
Multiply thru by 3w^2+34w-112 to get::
(3w-8) + (w+14) = -50
-------
4w + 6 = -50
4w = -56
w = -14
--------
Cheers,
Stan H.
==========